# CoAct-1: Computer-using Agents with Coding as Actions

### v2

@@ 4.2 Baselines
We compare CoAct-1 with OpenAI CUA 4o (OpenAI, 2025a), GTA-1 (Yang et al., 2025), UI-TARS (Qin et al., 2025), and other baselines listed in Table 1. These baselines represent the forefront of GUI-based task automation, each with a unique approach to interacting with and controlling the computer.
  • OpenAI CUA 4o (SOTA of accessible proprietary model). This agent uses vision and reasoning to interact with graphical user interfaces, controlling the mouse and keyboard to perform tasks. It is the technology behind services like Operator and ChatGPT agent.
  • GTA-1 (SOTA of grounder + planner) The GUI Test-time Scaling Agent (GTA-1) is a GUI agent that addresses the challenges of planning ambiguity and action grounding in high-resolution interfaces. To improve accuracy, this agent employs a "test-time scaling" strategy, where it generates multiple possible actions and utilizes a Judge model to select the optimal one. It also leverages the GRPO (Guo et al., 2025) to train a powerful grounder.
  • UI-TARS (SOTA of opensourced unified GUI agent). This is a self-contained agent that learns from a large dataset of GUI screenshots and computer-using trajectories. It utilizes a unified action model to execute human-like mouse and keyboard operations across various platforms, eliminating the need for external models.
@@ 4.3 Implementation Details
CoAct-1 settings We implement CoAct-1 using AG2 (Wu et al., 2023). We tried different backbones for the Programmer and Orchestrator, including OpenAI o3 (OpenAI, 2025b) and o4-mini (OpenAI, 2025b). For the GUI Operator, we use computer-use-preview, a vision-language action model finetuned by OpenAI for computer use, as the backbone model. We use the o4-mini as the summarizer for summarizing the conversation history between the Programmer and the Orchestrator. We set the max round $I$ for Programmer as 20, max step $K$ for GUI Operator as 25, and max round $J$ for Orchestrator as 15. Therefore, the number of system interactions, i.e, the number of steps, for CoAct-1 is upper bounded by 375.
Evaluation Protocol We evaluate our method with the rule-based evaluator provided by OSWorld. Internally, every evaluator is expressed as a Boolean expression built from 134 atomic, execution-based evaluators that the authors handcrafted for the benchmark. For a given task, the benchmark composes these atoms with logical AND / OR operators, so a “pass” might require, for instance, (file exported AND MD5 matches) AND (email sent == True).
@@ 4.4 Results
Table 1: Comparison of the state-of-the-art methods on the OSWorld (Xie et al., 2024) benchmark. We split the results by steps and show the approach type in the second column. A specialized model means that the model is trained specifically for computer use. We report the success rate (%) as the evaluation metric in the third column. We only include the verified results from https://os-world.github.io/.
      Agent Model       Approach Type       Success Rate
      15 steps
      o3 (OpenAI, 2025b)       General Model       9.10
      UI-TARS-72B-DPO (Qin et al., 2025)       Specialized Model       24.00
      UI-TARS-1.5-7B (Qin et al., 2025)       Specialized Model       25.70
      OpenAI CUA 4o (OpenAI, 2025b)       Specialized Model       26.00
      Jedi-7B w/ o3 (Xie et al., 2025b)       Agentic Framework       26.80
      Claude 3.7 Sonnet (Anthropic, 2025)       General Model       27.10
      Claude 4 Sonnet (Anthropic, 2025)       General Model       31.20
      Agent S2 w/ Gemini-2.5-Pro (Agashe et al., 2025)       Agentic Framework       34.64
      Agent S2.5 w/ o3 (Agashe et al., 2025)       Agentic Framework       39.00
      CoAct-1 w/ o3 & o4mini & OpenAI CUA 4o       Agentic Framework       39.81
      50 steps
      o3 (OpenAI, 2025b)       General Model       17.17
      UI-TARS-72B-DPO (Qin et al., 2025)       Specialized Model       25.80
      UI-TARS-1.5-7B (Qin et al., 2025)       Specialized Model       29.40
      OpenAI CUA 4o (OpenAI, 2025b)       Specialized Model       31.30
      Claude 3.7 Sonnet (Anthropic, 2025)       General Model       35.80
      Claude 4 Sonnet (Anthropic, 2025)       General Model       43.90
      Agent S2 w/ Gemini-2.5-Pro (Agashe et al., 2025)       Agentic Framework       45.76
      GTA-1-7B w/ o3 (Yang et al., 2025)       Agentic Framework       48.59
      Jedi-7B w/ o3 (Xie et al., 2025b)       Agentic Framework       50.60
      Agent S2.5 w/ o3 (Agashe et al., 2025)       Agentic Framework       54.20
      CoAct-1 w/ o3 & o4mini & OpenAI CUA 4o       Agentic Framework       56.39
      100 steps
      o3 (OpenAI, 2025b)       General Model       23.00
      UI-TARS-72B-DPO (Qin et al., 2025)       Specialized Model       27.10
      UI-TARS-1.5-7B (Qin et al., 2025)       Specialized Model       29.60
      OpenAI CUA 4o (OpenAI, 2025b)       Specialized Model       31.40
      Claude 3.7 Sonnet (Anthropic, 2025)       General Model       35.60
      Claude 4 Sonnet (Anthropic, 2025)       General Model       41.40
      Jedi-7B w/ o3 (Xie et al., 2025b)       Agentic Framework       51.00
      GTA-1-7B w/ o3 (Yang et al., 2025)       Agentic Framework       53.10
      Agent S2.5 w/ o3 (Agashe et al., 2025)       Agentic Framework       56.00
      CoAct-1 w/ o3 & o4mini & OpenAI CUA 4o       Agentic Framework       59.93
      100+ steps
      CoAct-1 w/ o3 & o4mini & OpenAI CUA 4o       Agentic Framework       60.76
Table 2: Per-task performance comparison on OSWorld tasks with 100 steps budget. The number beside each subtask is the total number of that subtask. It is obvious that CoAct-1 outperforms the previous SOTA methods in almost all subtasks and achieves the best overall average performance. Subtasks with the most performance gain are Multi-apps, Thunderbird, VLC, and OS, demonstrating the effectiveness of integrating a Programmer.
Methods (100 steps) Chrome (46) Calc (47) Impress (34) Writer (23) GIMP (25) VSCode (23) Multi Apps (101) Thunderbird (15) OS (24) VLC (17) Avg.
OpenAI CUA 4o 34.56 12.76 23.19 38.69 65.38 56.52 8.91 46.66 45.83 30.58 31.40
UI-TARS-1.5 (7b) 41.30 19.14 36.17 34.78 42.31 52.17 7.92 53.33 45.83 11.76 29.60
GTA-1 54.26 59.57 46.72 60.73 61.53 60.86 38.34 86.66 62.50 53.29 53.10
CoAct-1 52.09 70.21 50.27 73.91 65.38 78.26 47.88 66.67 75.00 66.07 59.93
Table 1 presents a comprehensive comparison of CoAct-1 with state-of-the-art models on the OSWorld benchmark, demonstrating that our agent establishes a new standard for performance. CoAct-1 achieves a top success rate of 60.76% in the 100+ step category. The results show that CoAct-1 consistently outperforms other leading methods across various maximum step allowances. For instance, at a 100-step budget, CoAct-1 reaches a 59.93% success rate, which is a significant improvement over other agentic frameworks like Agent S2.5 w/ o3 (56.00%) and GTA-1-7B w/ o3 (53.10%). This robust performance is also evident at the 50-step limit, where CoAct-1 achieves 56.39%, surpassing all other agents.
On the other hand, as shown in Table 2, CoAct-1 achieves the highest overall average success rate of 59.93%. This represents a significant margin of improvement over other leading agents, including GTA-1 (53.10%), OpenAI CUA 4o (31.40%), and UI-TARS-1.5 (7b) (29.60%). The advantages of our hybrid approach are most evident in task categories that benefit from programmatic control. In the "Multi Apps" category, CoAct-1’s performance of 47.88% substantially exceeds that of GTA-1 (38.34%). Similarly, major performance gains are observed in application-specific and OS-level tasks. For VLC, CoAct-1 scores 71.96% compared to GTA-1’s 53.29% ; and for general OS tasks, CoAct-1 reaches 75.00%. CoAct-1 also demonstrates superior performance in other applications like LibreOffice Calc (70.21%) and Writer (73.91%). These results underscore the effectiveness of integrating a Programmer agent, which enables CoAct-1 to handle a more diverse and complex range of tasks by augmenting GUI manipulation with the precision of direct coding actions.
@@ 4.5 Discussion
(a) Average steps for a successful task. CoAct-1 can solve tasks with lower average steps than previous SOTA agentic frameworks. On the other hand, CoAct-1 higher steps than OpenAI CUA 4o but also higher accuracy.
(b) GUI/code ratio split by total action of tasks. Coding helps reduce the total action steps.
(c) This chart illustrates the domains where coding actions are most frequently applied.
(d) Error rate distribution split by total action of tasks. The error rate is positive related to total actions.
@@
The analysis of CoAct-1’s operational efficiency, illustrated in Figure 3, reveals that our hybrid approach is substantially more efficient than leading GUI-only agents. This efficiency is a key factor in its improved success rate. As shown in 3(a), CoAct-1 solves tasks with an average of 10.15 steps. This represents a significant improvement over other high-performing agents like GTA-1, which requires 15.22 steps, and UI-TARS, which needs 14.90 steps on average. While OpenAI CUA 4o averages fewer steps (6.14) , its overall success rate is much lower compared to CoAct-1’s (31.40% v.s. 59.93% on 100 steps). This indicates that CoAct-1’s efficiency is coupled with greater effectiveness. The source of this efficiency lies in the strategic use of coding actions. 3(b) supports this by showing that coding actions help keep the total steps per task relatively low. This efficiency is crucial for robust performance. 3(c) shows that coding is particularly beneficial in complex domains like "LibreOffice Calc", "Multi-apps" and direct OS interactions, where a large proportion of tasks are solved with code. A single script can replace a long and error-prone sequence of GUI clicks, streamlining the workflow. 3(d) illustrates a clear trend: tasks that require more actions are more likely to fail. By reducing the total number of steps, the hybrid approach not only accelerates task completion but also minimizes the opportunities for error. The ability to dynamically select the most appropriate action—either a direct coding command or a GUI interaction—is fundamental to the enhanced efficiency and reliability of CoAct-1.
@@
CoAct-1 OpenAI CUA 4o o4-mini o4-mini 47.41
o3 o3 53.12
@@
We investigated the impact of different backbone models on the performance of CoAct-1, with results presented in Table 3. The study reveals that the choice of models for the Orchestrator, Programmer, and GUI Operator significantly influences overall effectiveness. When a uniform backbone, o4-mini, was used for both Orchestrator and Programmer, the system achieved a performance score of 47.41%. Performance improved to 53.12% when the o3 model serving as both the Orchestrator and Programmer. The highest performance of 60.76% was achieved by retaining OpenAI CUA 4o for the GUI Operator and o3 for the Orchestrator, while upgrading the Programmer to the o4-mini model. This highlights that the system benefits from a powerful, vision-centric model for GUI operations and that enhancing the capability of the Programmer agent yields significant performance gains.
4.6 Case Study

### v3

@@ 4.2 Baselines
We compare CoAct-1 with two categories of computer-using agents, the end-to-end models and agentic methods. These baselines represent the forefront of GUI-based task automation.
End-to-end model
An end-to-end model takes user instructions and OS screenshots as input and outputs corresponding actions by pure inference without any agentic workflow. OpenAI o3 (OpenAI, 2025b) is a cutting-edge reasoning model from OpenAI that excels at multi-step problem-solving and versatile, context-aware assistance. OpenAI CUA 4o (OpenAI, 2025a) uses vision and reasoning to interact with graphical user interfaces, controlling the mouse and keyboard to perform tasks. It is the technology behind services like Operator and ChatGPT agent. UI-TARS (Qin et al., 2025) UI-TARS introduces a fully end-to-end, screenshot-only native GUI agent model that unifies perception, reasoning, memory, and action. OpenCUA(32B) (Wang et al., 2025b) introduces a fully open-source framework, including a scalable data collection tool, the first large-scale multi-OS computer-use task dataset, a reflective chain-of-thought reasoning pipeline, and strong vision-language agent models.
Agentic Method
The agentic method encompasses single- and multi-agent systems with diverse structures, such as planner-grounder, planner-multigrounder, etc. These baselines primarily focus on enhancing the ground-level ability of a language model to improve computer-based performance. Jedi-7B w/ o3 (Xie et al., 2025a) We refer the term Jedi to a Qwen2.5-VL trained on the Jedi dataset. The author plugs Jedi into an agent stack to translate high-level plans into pixel-perfect GUI actions, achieving large gains on OSWorld, WindowsAgentArena, and multiple grounding benchmarks. GTA-1 w/ o3 (Yang et al., 2025) The GUI Test-time Scaling Agent (GTA-1) is a GUI agent that addresses the challenges of planning ambiguity and action grounding in high-resolution interfaces. To improve accuracy, this agent employs a "test-time scaling" strategy, where it generates multiple possible actions and utilizes a Judge model to select the optimal one. It also leverages the GRPO to train a powerful grounder. Agent S2.5 w/ o3 (Agashe et al., 2025) Agent S2.5 is a compositional planner-multigrounder framework in which a planner generates high-level subgoals, multiple grounders executes them while delegating GUI-element localization to visual, textual, and structural experts via a Mixture-of-Grounding, and both levels proactively replan after every subgoal to remain robust to changing screens.
Besides the above powerful baselines, we also add Agent S (Agashe et al., 2024) and NAVI (Bonatti et al., 2024) as baselines for WindowsAgentArena.
@@ 4.3 Implementation Details
CoAct-1 settings We implement CoAct-1 using AG2 (Wu et al., 2023). In CoAct-1, we adopt OpenAI o3 for Orchestrator and OpenAI o4-mini for Programmer. For the GUI Operator, we use OpenAI computer-use-preview, a vision-language action model finetuned by OpenAI for computer use, as the backbone model. We use the o4-mini as the summarizer for summarizing the conversation history between the Programmer and the Orchestrator. We set the maximum round $I$ for the Programmer to 20, the maximum step $K$ for the GUI Operator to 25, and the maximum round $J$ for the Orchestrator to 15. Therefore, the number of system interactions, i.e, the number of steps, for CoAct-1 is upper bounded to 375 (but in all cases, as shown in 2(d), CoAct-1 will early stop before 150 steps). More details are in Appendix C.
Evaluation Protocol We evaluate our method with the rule-based evaluator provided by OSWorld and WindowsAgentArena. Internally, every evaluator is expressed as a Boolean expression built from 134 atomic, execution‑based evaluators that the authors handcrafted for the benchmark. For a given task, the benchmark composes these atoms with logical AND / OR operators, so a “pass” might require, for instance, (file exported AND MD5 matches) AND (email sent == True).
@@ 4.4 Results
Table 1: Comparison of the state-of-the-art methods on the OSWorld (Xie et al., 2024) verified benchmark. We split the results by steps and show the approach type in the second column. Office tasks include tasks from LibreOffice Calc, LibreOffice Impress, and LibreOffice Writer. Daily tasks include tasks from Chrome, Thunderbird, and VLC. Professional includes tasks from GIMP and VSCode. We report the success rate (%) as the evaluation metric for each type of task, and mark the best result of each step budget in bold and the best result overall step budgets in underlined bold.
Agent Model Office (104 tasks) Daily (78 tasks) Professional (48 tasks) OS (24 tasks) Multiple Apps (101 tasks) Avg.
[…]
OpenAI o3 1.45 8.02 12.29 37.50 11.82 9.09
UI-TARS-1.5 (7B) 27.19 27.99 61.45 34.78 5.38 25.76
OpenAI CUA 4o 22.17 37.65 41.22 45.83 10.75 26.01
OpenCUA 26.69 33.25 51.57 43.48 10.41 28.12
Agent S2.5 w/ o3 42.85 44.61 57.10 70.83 17.82 38.98
Jedi-7B w/ o3 45.84 57.49 60.95 50.00 20.43 42.37
[…]
OpenAI o3 11.50 19.78 30.10 37.50 11.82 17.17
UI-TARS-1.5 (7B) 26.51 31.41 48.91 25.00 9.77 25.08
OpenAI CUA 4o 23.56 38.43 52.09 70.83 15.86 31.19
OpenCUA 30.06 42.31 58.28 47.83 16.79 33.76
GTA-1-7B w/ o3 48.58 52.31 77.84 58.33 37.05 48.59
Jedi-7B w/ o3 50.10 65.25 68.65 54.17 34.97 50.65
Agent S2.5 w/ o3 52.81 55.80 75.42 75.00 39.53 54.21
[…]
OpenAI o3 17.23 26.29 38.79 62.50 16.53 23.00
UI-TARS-1.5 (7B) 25.01 31.07 46.99 29.17 8.80 25.41
OpenAI CUA 4o 25.04 39.19 55.43 58.33 18.48 31.38
OpenCUA 30.06 38.89 60.70 52.17 18.10 33.84
Jedi-7B w/ o3 47.89 64.37 75.92 50.00 35.27 50.98
GTA-1-7B w/ o3 55.68 64.74 61.20 62.50 38.34 53.07
Agent S2.5 w/ o3 54.23 55.80 75.42 75.00 44.06 55.98
[…]
Table 2: Comparison of the state-of-the-art methods on the WindowsAgentArena (Bonatti et al., 2024). We split the results by steps and show the approach type in the second column. Office tasks include tasks from LibreOffice Calc and LibreOffice Writer. Web tasks include tasks from Chrome, and Microsoft Edge. Windows System includes tasks from settings and File Explorer. Windows System includes tasks from Settings and File Explorer. Windows Utils includes tasks from Clock, Windows Calculator, Notepad, and Microsoft Paint. We report the success rate (%) as the evaluation metric for each type of task and mark the best result in bold.
Method Office(43 tasks) Web(30 tasks) Windows System(24 tasks) VSCode(24 tasks) VLC(21 tasks) Windows Utils(12 tasks) Avg.
Agent S 0.0 13.3 45.8 29.2 19.1 22.2 18.2
NAVI 0.0 27.3 33.3 27.3 30.3 8.3 19.5
Agent S2 7.0 16.4 54.2 62.5 28.6 33.3 29.8
[…]
Our experimental results, detailed in subsection 4.4 and Table 2, unequivocally establish CoAct-1 as the new state-of-the-art across two challenging, real-world computer operation benchmarks. The findings validate our core hypothesis: integrating programmatic actions alongside traditional GUI manipulation provides a more robust, efficient, and generalizable paradigm for computer automation.
Performance on OSWorld
On the comprehensive OSWorld benchmark (subsection 4.4), CoAct-1 demonstrates superior performance and efficiency. It achieves a final success rate of 60.76% within the 150-step limit, creating a significant margin over the strongest contemporary agentic frameworks, including Agent S2.5 w/ o3 (55.98%) and GTA-1 w/ o3 (53.07%). The strength of our hybrid architecture is not only in its peak performance but also in its consistency across different task complexities. At the 100-step mark, CoAct-1 already leads with a 59.93% success rate, surpassing the final scores of all other baselines. The advantages of our approach are particularly pronounced in categories where programmatic control is most effective. CoAct-1 achieves top-tier performance in tasks requiring complex data or file system interactions, scoring 64.80% in Office, 75.00% in OS, and 47.87% in Multiple Apps. This exceptional performance in domains that are historically brittle for pure-GUI agents underscores the efficacy of delegating backend operations to the Programmer agent, which can execute precise and reliable scripts for tasks involving spreadsheets, file manipulation, and cross-application data flows.
Performance on WindowsAgentArena
We further evaluated CoAct-1 on the WindowsAgentArena benchmark (Table 2). The results show that our framework successfully transfers its capabilities to a different operating system, again achieving state-of-the-art performance by a substantial margin. CoAct-1 attains an overall success rate of 52.5%, which is a remarkable improvement over prior leading methods like Agent S2 (29.8%) and NAVI (19.5%). The performance breakdown on WindowsAgentArena further reinforces our central claim. CoAct-1 shows commanding strength in system-level and utility-based tasks, achieving standout scores of 83.3% in Windows System and 77.7% in Windows Utils. These categories, which involve interacting with file explorers, system settings, and other native utilities, are ideally suited for the script-based actions of the Programmer. Furthermore, CoAct-1’s performance demonstrates clear and effective scaling with an increased step budget, rising from 21.4% at 15 steps to 52.5% at 100 steps, highlighting its capacity to solve more complex, long-horizon problems.
In summary, the consistent, state-of-the-art performance across two distinct and challenging benchmarks confirms that CoAct-1’s hybrid agentic architecture represents a significant advancement toward creating more capable and reliable autonomous agents for general computer use.
@@ 4.5 Discussion
(a) Average steps per passed task with 100 step budget, showing CoAct-1 is significantly more efficient than other SOTA agentic frameworks like GTA-1.
(b) Distribution of tasks by total step count, illustrating the ratio of coding to GUI actions and showing that coding helps reduce the total action steps.
(c) Breakdown of passed tasks by application domain, highlighting that coding actions are most frequently applied in complex domains like LibreOffice Calc, Multi-Apps, and direct OS interaction.
(d) Pass/fail ratio split by total steps, demonstrating that the failure rate is positively correlated with the number of actions required.
@@
The analysis of CoAct-1’s operational efficiency, illustrated in Figure 3, reveals that our hybrid approach is substantially more efficient than leading GUI-only agents. This efficiency is a key factor in its improved success rate. As shown in 2(a), CoAct-1 solves tasks with an average of 10.15 steps. This represents a significant improvement over other high-performing agents like GTA-1, which requires 15.22 steps, and UI-TARS, which needs 14.90 steps on average. While OpenAI CUA 4o averages fewer steps (6.14), its overall success rate is much lower compared to CoAct-1’s (31.40% v.s. 59.93% on 100 steps). This indicates that CoAct-1’s efficiency is coupled with greater effectiveness. The source of this efficiency lies in the strategic use of coding actions. 2(b) supports this by showing that coding actions help keep the total steps per task relatively low. This efficiency is crucial for robust performance. 2(c) shows that coding is particularly beneficial in complex domains like "LibreOffice Calc", "Multi-apps", and direct OS interactions, where a large proportion of tasks are solved with code. A single script can replace a long and error-prone sequence of GUI clicks, streamlining the workflow. 2(d) illustrates a clear trend: tasks that require more actions are more likely to fail. By reducing the total number of steps, the hybrid approach not only accelerates task completion but also minimizes the opportunities for error. The ability to dynamically select the most appropriate action—either a direct coding command or a GUI interaction—is fundamental to the enhanced efficiency and reliability of CoAct-1.
@@
OpenAI CUA 4o o4-mini o4-mini 43.43
o3 o3 58.72
@@
We investigated the impact of backbone model selection for each agentic component of CoAct-1 on the OSWorld benchmark, with results presented in Table 3. Our analysis reveals that the overall system performance is highly sensitive to the reasoning and instruction-following capabilities of the models chosen for the Orchestrator and Programmer roles. When utilizing o4-mini for both the Orchestrator and Programmer, alongside the OpenAI CUA 4o as the GUI Operator, the system achieved a performance of 43.43%. A significant performance enhancement to 58.72% is observed when a more powerful model, o3, is used for both the Orchestrator and Programmer. This underscores the critical role of a sophisticated high-level planner and a capable code generator in the system’s success. The highest performance of 60.76% was achieved with a heterogeneous configuration: employing o3 for the Orchestrator, o4-mini for the Programmer, and retaining OpenAI CUA 4o for the GUI Operator. This configuration suggests an optimal balance, leveraging the powerful reasoning of o3 for task decomposition and delegation, while benefiting from the specialized capabilities of o4-mini for code generation. These results highlight that enhancing the capabilities of the Orchestrator and Programmer yields the most substantial performance gains, validating our modular design and demonstrating the benefits of strategically allocating powerful models to roles with high reasoning demands.
Pure GUI action VS pure coding action
Table 4: Ablation study on the performance of CoAct-1’s components. We compare the full hybrid system against agents restricted to using only the Programmer (pure coding) and only the GUI Operator (pure GUI). The results highlight that the integrated CoAct-1 system significantly outperforms either single-modality agent, demonstrating the effectiveness of its hybrid approach.
CoAct-1 Office Daily Professional OS Multiple Apps Avg. Avg. Steps
w/ Programmer w/ GUI Operator
40.88 16.17 53.06 62.50 29.63 35.73 1.14
43.50 58.80 69.38 79.16 35.68 50.68 11.20
64.80 66.51 71.82 75.00 47.87 60.76 10.15
To isolate each agent modality and validate our hybrid design, we conducted an ablation study comparing the full CoAct-1 system with agents restricted to a single action type. Results in Table 4 show the combined approach is superior. A Programmer-only agent (pure coding) achieved 35.73% success, highlighting that many tasks require GUI interaction beyond scripting. It was highly efficient, averaging just 1.14 steps per success, confirming the directness of programmatic actions. A GUI Operator–only agent (pure GUI) achieved a higher 50.68% success rate, handling more task types, but required 11.20 steps per task. The full CoAct-1 model, integrating both modalities, achieved 60.76% success with 10.15 steps on average, demonstrating the synergy of our architecture: the Orchestrator exploits the Programmer’s efficiency for backend tasks and the GUI Operator’s versatility for visual navigation, yielding a robust system.
5 Conclusions
In this work, we introduced CoAct-1, a novel multi-agent system designed to address the inherent inefficiency and brittleness of agents that rely exclusively on GUI manipulation. Our multi-agent system features an Orchestrator that dynamically delegates subtasks to a GUI Operator or a Programmer. Our extensive evaluation on the OSWorld and WindowsAgentArena benchmark confirms the effectiveness of this approach. CoAct-1 achieves a new state-of-the-art success rate of 60.76% on OSWorld and 52.5% on WindowsAgentArena, significantly outperforming previous leading methods. The performance gains were particularly pronounced in categories involving OS-level interactions, multi-application workflows, and other tasks where the Programmer agent could leverage direct programmatic execution.
Appendix A LLM Usage Statement
We used a large language model (OpenAI’s o3 and GPT-5) as a general-purpose writing assistance tool. Its role was limited to sentence- and paragraph-level polishing, including improving clarity, grammar, and flow. The authors developed all research ideas, analyses, results, and conclusions. The model did not generate new content, perform a literature review, or contribute to the conceptual framing of the paper.
Appendix B Ethic Statement
Our research aims to advance computer automation for beneficial and productive purposes, but we acknowledge the potential for dual-use and associated risks.
Security and Misuse: An agent with the ability to execute code (Python or Bash) and manipulate a GUI could be leveraged for malicious activities if not properly constrained. To mitigate this risk during our research, all experiments were conducted within secure, isolated, and virtualized benchmark environments (OSWorld and WindowsAgentArena). This ensures that the agent’s actions are sandboxed and cannot affect real-world systems or data. We advocate that any future deployment of such agents in live environments must incorporate robust security protocols, strict permission controls, and mechanisms to prevent the execution of harmful code.
Data Privacy: The agent’s operation relies on observing the screen via screenshots, which in a real-world scenario could contain sensitive or personal information. In this work, we use only the data provided within the benchmark tasks, which do not involve real user data. For any future applications, it is imperative to implement strict data handling policies and privacy-preserving techniques to protect user confidentiality.
Appendix C Reproducibility Statement
Reproduction of CoAct-1 requires an accurate use of specific OpenAI models and prompts. Please review the following details to ensure the accurate performance of our work.
C.1 Model Usage and Environment Setting
Model Usage
In this work, we use o3-2025-04-16 for Orchestrator, o4-mini-2025-04-16 for Programmer, and computer-use-preview-2025-03-11 for GUI operator. Any open-sourced model can also be adopted by CoAct-1 if it meets the following requirements:
  • For Orchestrator The Orchestrator requires multi-modality input (image and text) to process all screenshots from the OS for planning. It also requires a strong reasoning ability across different modalities (See Table 3, reasoning ability will largely affect the final performance).
  • For GUI Operator The GUI Operator can be replaced with any open-sourced vision language action models (VLA) like UI-TARS (Qin et al., 2025) or OpenCUA (Wang et al., 2025b), or planner-grounder approaches like GTA-1 (Yang et al., 2025) and Agent S2 (Agashe et al., 2025). CoAct-1 require the GUI Operator to have the following two abilities: (1) instruction following ability out of the grounding task (like when to terminate), and (2) computer use ability, including clicking, dragging, typing, and hotkeys. Unfortunately, we have not been allocated enough GPU resources for testing these models’ performance on CoAct-1 in this work, and you can expect a reduction in average steps and a performance improvement when switching to a stronger model for the GUI Operator.
  • For Programmer Language model for Programmer should have a strong ability to write Python and Bash (or Powershell in Windows) for solving computing problems. Note that we didn’t provide any API or SDK for programmers; however, this can be an extension to your approach for better performance on app-specific tasks, such as Chrome and Thunderbird.
Environment
The OS environment is also a significant challenge when evaluating computer-using agents. To reproduce our experiment results, all input screenshots to the Orchestrator should accurately reflect the well-initialized system state. Therefore, we recommend waiting 60 seconds after the VM starts before capturing the OS screenshot to ensure the screenshot includes all necessary information for Orchestrator to plan.
C.2 Prompt Design for CoAct-1
The performance of any modern large language model agent system can be largely affected by its prompt or template design (He et al., 2024). In CoAct-1, despite the rules for general behavior like chain-of-thought and verification, we also design rules based on the model limitations. Specifically, we notice that the GUI Operator has a significant hallucination rate when performing self-checks, while the Programmer’s file modifications do not reflect in the OS, so the Orchestrator cannot capture and plan for the next step. We mitigate these limitations in our prompt design by allowing Orchestrator to check the result independently and reload the file modified by the Programmer. We put the prompt used for OSWorld and WindowsAgentArena in Table 5, Table 6, Table 7, Table 8, and Table 9.
Appendix D When&Why CoAct-1 Fails? Insight for Future Works